OpenStudy (mellamokatie):

a 154 -lb person burns 430 calories per hous riding an exercize bicycle at a rate of 15 mi/hr. write a fucntion rule to represent the total calories burned over time by that person. explain how the information in the problem relates to the function

OpenStudy (mellamokatie):

@sweetburger @jim_thompson5910

jimthompson5910 (jim_thompson5910):

There's a lot of useless info in this problem meant to throw you off. The only thing you need is the fact that you burn 430 calories per hour if x is the number of hours and y is the amount of calories burned, then y = 430*x

jimthompson5910 (jim_thompson5910):

In function notation, you replace the y with f(x) to get f(x) = 430*x

jimthompson5910 (jim_thompson5910):

Let's say x = 1 f(x) = 430*x f(1) = 430*1 ... replace every x with 1 f(1) = 430 so in 1 hour of exercise, you burn off 430 calories as expected

OpenStudy (mellamokatie):

thanks so much i Really appreciate this! i have a couple more could you help on 2 more? please

jimthompson5910 (jim_thompson5910):

`i have a couple more could you help on 2 more? ` sure I can help

jimthompson5910 (jim_thompson5910):

please make a new post for the new question

OpenStudy (mellamokatie):

one second!